the digit in the tens place of a two digit number is six more than twice the digit in the ones place. If the two digits were reversed, the new number will be 63 less than the original number. what is the original number?

Answer by Boreal(15235)

Let x = the digit in the units column
Then 2x+6 is the digit in the tens column twice the number plus 6
One number is 10(2x+6) +x The 10 is for the tens column, and +x is the units.
Think of 24 as being 10*2 +4 Same approach here.
Other number is 10x +(2x+6)
The first is 63 more than the second.
Therefore 10(2x+6) +x-63=10x +2x+6 The minus 63 means the first number is 63 more than the second, so we have to subtract 63 to make them equal.
Distribute
20x +60+x -63=12x +6
Collect terms
21x +60-63=12x +6
Solve
9x -3 =6
9x=9
x=1
2x+6=8
81 is the first number, 18 is the second number. They are 63 apart.